I go 6 blocks north, 3 blocks east, 2 blocks south, 4 blocks west, and finally, another 4 blocks south. What distance have I traveled? What is my total displacement?
1
Expert's answer
2012-09-25T10:45:59-0400
Distance D = 6 + 3 + 2 + 4 + 4 = 19 blocks.
Assuming north and east
directions as positive, getting:
N = 6 - 2 - 4 = 0
E = 3 - 4 = -1
So,
the total displacement L = sqrt( N^2 + E^2 ) = sqrt(0 + 1) = 1 block.